The length of a rectangle is five times its width.
If the area of the rectangle is 245 ft, find its perimeter.

Answer by rm29924(97)

let the width of the rectangle = w
the length of the rectangle = 5w
to get the area of a rectangle you multiply the length times the width
5w(w)=245 ft
5w^2=245 ft
w^2=49
w=7
therefore the width = 7 ft and the length = 35 ft
the formula to find the perimeter is 2w + 2l
2(7) + 2(35)
14 + 70 = 84
The perimeter is 84 ft.